9, Award: 0 out of 10.00 points Which quantity represents the smallest amount of radioactivity? 1 x 1013 Ci O 1x 104 mOi O 1x1010 uci References Multiple Choice Difficulity: Hard
Answer #1

1 Bq = 2.7 x 10 -11 Ci

1mCi = 1 x 10-3 Ci, thus, 1 x 10 4 mCi = (1 x 10 4) 1 x 10-3 Ci = 10 Ci

1\mu Ci = 1 x 10 -6 Ci . thus,  10 \mu Ci = 10 x 1 x 10 -6 Ci = 1 x 10 -5 Ci

Thus, the smallest amount of radioactivity is represented by 1 Bq (= 2.7 x 10 -11 Ci )